mr c can type 66 words in 1.5 minutes. at the same rate how many words can he type in 3.5 minutes?
Answer by Kalmetam(43)

yeah let me try this in an easier way..
so he can type 66 in 1.5 minutes.. ok
Well if we add 2 to 1.5, we get 3.5, so we need to use a simple fraction of 1.5 that the denominator is a factor of 66, then we can multiply that to the amount needed to get 3.5
so lets see...
1/3 of 1.5 = x
Well 1/3 of 15 is 5, so this is .5
so..
x=.5
so lets see, we divided 1.5 by 3, now we need to divide 66 by 3.
66/3=y
so...
66/3=22

so now we have 22 words per 0.5 minutes.
now we need to know how many times that goes into 3.5
3.5/x=z
3.5/.5=z
z=7
and now you need to multiply the OTHER SIDE!
7y=u
7*22=u
u=154
he can type 154 words in 3.5 minutes..